OpenStudy (anonymous):

13 times 6 divided by 3 times 2 divided by 13.

OpenStudy (ja1):

Ok do order of operations

OpenStudy (ja1):

13 * 6 = ?

OpenStudy (ja1):

13 * 6 = 78 78/3 = 26 26 * 2 = 52 52 / 13 = 4

OpenStudy (anonymous):

P - paranthesis E - exponents M - multiplication D - division A - addition S - subtraction

OpenStudy (ja1):

@grayp do you understand or do you want me to explain?

OpenStudy (anonymous):

\[13 \times \frac{ 6 }{ 3 } \times \frac{ 2 }{ 13 } = 4\]

OpenStudy (anonymous):

the answer is 9

OpenStudy (ja1):

What? How?

OpenStudy (anonymous):

how did you get 9?

OpenStudy (ja1):

My question exactly

OpenStudy (anonymous):

I meant to say four, my bad. thanks to both of you though.

OpenStudy (ja1):

lol no prob glad to help :)